OpenStudy (anonymous):

Please help!!!!! I will give a medal!!!!1 For the following system, if you isolated x in the first equation to use the Substitution Method, what expression would you substitute into the second equation? -x - 2y = -4 3x + y = 12 -2y - 4 2y - 4 2y + 4 -2y + 4

OpenStudy (anonymous):

x = -2y +4 just make x subject of formula in first equation :)

OpenStudy (anonymous):

See, im new to this. How exactly would i do that.

OpenStudy (anonymous):

Do i take away x and make it -2y + 4 +x

OpenStudy (anonymous):

\[- x -2y = -4\] \[-x = -4 + 2y\] \[( \times -1) : -x \times -1 = -4 \times -1 + 2y \times -1\] \[x = 4 -2y\] which is same as : \[x = -2y +4\]

OpenStudy (anonymous):

here is what i did to get x as Subject of formula if you dont understand some part just ask :)

OpenStudy (anonymous):

I don't understand any of it.

OpenStudy (anonymous):

ok you have equation number 1 : - x - 2y = -4 now you need an equation whereby you need to substitute the x value of the equation number 1 into equation number 2

OpenStudy (anonymous):

positive x- value : x = -2y +4

OpenStudy (anonymous):

Oh okay, thank you!